OpenStudy (anonymous):

14+6a-8=18 i think the answer is 2 but i dont know how to get it

OpenStudy (anonymous):

I'll help :)

OpenStudy (anonymous):

Please

OpenStudy (anonymous):

14+6a-8=18 14+6a=18+8 14+6a=26 14-14+6a=26-14 6a=12 a=12/6 a=2

OpenStudy (anonymous):

subtract 14-6 you should get 6! so now your equation is 6a+6=18 subtract 6 from both sides. now you have 6a=12 divide both sides by 6 to get a which is 2

OpenStudy (anonymous):

your answer was correct,i love your avatar by the way

OpenStudy (anonymous):

Im in 8th grade doing 9th grade math so i skipped the regular 8th grade math. THANK YOU!

OpenStudy (anonymous):

@Tris

OpenStudy (anonymous):

Great!I am in 8th grade too

OpenStudy (anonymous):

Yea im in Pre-AP so i got Algebra 1

OpenStudy (anonymous):

wow thats pretty impressive i think...so your christian?I should get a nice avatar like that,my annoynymus bird is killing me

OpenStudy (anonymous):

Yes Faithful and Loyal to my One and Only God